OpenStudy (anonymous):

Rob was negotiating a new contract with is team. He wanted $700,000 for the year and an additional $1800 for each game. His team was offering him $600,000 for the year, and 5000 for every game started. How many starts would he need to make in order to earn more with the teams offer? PLEASE HELP ME!

OpenStudy (deoxna):

Make two equations where the total money payed is y and x is the number of games played. y=700000+1800x y=600000+5000x Plot both lines in a calculator and eventually the second one will move over the first, The x value of the intersection point is your answer.

OpenStudy (anonymous):

Or both equations equal to each other and solve for x.